# Not All Attention Heads Are What You Need: Refining CLIP’s Image Representation with Attention Ablation

### v1

@@ 6 Conclusion
To further understand the role of the attention mechanism in CLIP, we conduct a comprehensive analysis of all attention heads in its image encoder. Based on the hypothesis that certain heads may negatively impact downstream performance, we propose AAT, a simple yet effective method that refines output representations by systematically ablating detrimental heads through attention weight manipulation. To identify which heads to ablate, we introduce two alternative strategies: GA and BP. AAT-GA is well-suited for scenarios with limited floating-point computation, such as inference-friendly edge devices, while AAT-BP offers greater optimization efficiency when computational resources are less constrained. Extensive experiments demonstrate that AAT is a versatile and powerful technique for enhancing CLIP-family models in diverse downstream tasks. Ultimately, we hope this study contributes to a deeper understanding of CLIP and paves the way for more efficient use of vision-language models.
Appendix A Benchmark Details and Evaluation Metrics
A.1 MS COCO
@@ A.4 COCO-CN
A.5 Flickr30k-CNA
Flickr30k-CNA is a Chinese-translated version of Flickr30k, with text generated by professional English and Chinese linguists. We process and evaluate this dataset in the same manner as Flickr30k.
A.6 ImageNet-1k
@@
A.7 Cola
Cola [29] is a compositional text-to-image retrieval benchmark designed to evaluate a model’s ability to localize and compose objects with their corresponding attributes. It includes both real-world images and synthetic 3D-rendered scenes. The task requires retrieving images that match the correct configuration of objects and attributes while rejecting distractors that contain the right components in incorrect arrangements. Cola consists of approximately 1.2k composed queries involving 168 objects and 197 attributes, distributed across roughly 30K images. It features two types of queries: single-object compositions (sourced from GQA, CLEVR, and PACO) and multi-object queries. Following the original paper, mean average precision (mAP) is used for evaluating single-object subsets, while mean accuracy is used for the multi-object subset.
Appendix B Detailed Impact of Individual Attention Heads
In the main content of the paper, we identify all negatively impactful heads—those that individually yield higher retrieval performance than the vanilla baseline when ablated—and jointly ablate them in a simple strategy referred to as naive joint head ablation (naive h-a). We extend this experiment across different model sizes for both text-to-image and image-to-text retrieval on COCO-CN. As shown in Table 12, even this straightforward method surpasses the baseline by 1.4$\sim$1.6on the test set and 1.5$\sim$2.6on the all set for text-to-image retrieval, and by 0.4$\sim$1.2on the test set and 1.8$\sim$3.7on the all set for image-to-text retrieval.
@@ Appendix B Detailed Impact of Individual Attention Heads
Appendix C Analysis of the Ablated Heads
C.1 Layer Distribution
@@
01234567891011005555101010103333777755554444333333330033334444555566661111222244444444333333332222555533331111444477773333layersnumCLIPChinese-CLIP    (a) ViT-B-based AAT 0246810121416182022005555101010101515151555558888666644445555555544445555555577775555111111114444555566667777555500121212121010101088886666666699997777666655553333333388885555555555551010101066668888666666665555777766666666777766661010101066664444layersnumCLIPChinese-CLIP        (b) ViT-L-based AAT 024681012141618202224262830005555101010101515151577771010101088888888555555554444999955557777555522227777555511118888999977777777777799997777444488888888333355554444999955559999888810101010101010106666999912121212555588884444444499997777666611118888333344448888555555556666666677778888888866664444666677774444444488886666layersnumCLIPChinese-CLIP                (c) ViT-H-based AAT
@@ C.2 Statistics
Appendix D Computational Cost Analysis
We optimize AAT-GA by eliminating redundant overhead, as described in Section 5.4 of the paper. Detailed performance and runtime comparisons under different hyper-parameter configurations are provided in Tables 14 and 15.
@@ Appendix D Computational Cost Analysis
Table 14: Mean-R (mR) across different CLIP model sizes for text-to-image retrieval on the MS COCO test set, along with the corresponding GA optimization runtime (T, in minutes). For clarity, “gen” refers to the number of generations, and “psize” denotes the population size.
@@
Appendix E Evaluation on Compositional Retrieval
@@
Appendix F Comparison with SFT
To further validate the effectiveness and practicality of AAT, we compare it against conventional supervised finetuning (SFT). We conduct this comparison using Chinese-CLIP, where AAT demonstrates more pronounced improvements than on the English counterpart, making performance differences easier to observe. Since AAT are optimized using 1k image-text pairs from COCO-CN, we train the vanilla Chinese-CLIP on the same dataset using the contrastive supervision objective employed in CLIP, following the recommended hyper-parameters from the OpenCLIP codebase [15]. Specifically, we use a base learning rate of 5e-5 with a cosine decay schedule, the AdamW optimizer, a batch size of 256, and a weight decay of 1e-3. In experiments, we vary the number of training epochs from 1 to 64 to explore optimal performance across different model sizes. We evaluate the finetuned models on both the in-domain COCO-CN test set and the cross-domain Flickr30k-CNA test set for text-to-image retrieval. The results are presented in Figure 5.
SFT exhibits varying mean-R on COCO-CN across model sizes. It slightly outperforms AAT on ViT-B ($+$0.2%), achieves a larger gain on ViT-L ($+$2.2%), and performs comparably on ViT-H (--0.1%). However, its results are highly sensitive to the number of training epochs, likely due to overfitting or convergence to local minima under limited data. This instability is consistent across repeated trials, highlighting the challenges of SFT in data-scarce settings.
Appendix G Visualization of Samples for AAT Optimization and Its Failure Cases
For the convenience of readers, we provide a visualization of the validation set $\mathcal{D}$ used for AAT optimization in Figure 6. As discussed in Appendix E and Section 5.2.1 of the paper, AAT fails on Cola-CLEVR and three non-natural subsets from ReCoS. Representative examples from these tasks are shown in Figure 7 and Figure 8, respectively. As illustrated, there exists a significant domain gap between the test images and those in $\mathcal{D}$. Moreover, certain tasks require models to perform high-level comprehension and reasoning, which exceeds the representational capacity of CLIP.
@@ Appendix G Visualization of Samples for AAT Optimization and Its Failure Cases
Figure 8: Visualization of non-natural images from three ReCoS-v1 subsets: code reasoning (CR), numerical calculation (NC), and text translation (TT).

### v2

@@ 6 Conclusion
To better understand CLIP’s attention mechanism, we conduct comprehensive analysis of attention heads in its image encoder. Observing heterogeneous head impacts, we propose AAT, a simple but effective method that refines representations by systematically ablating detrimental heads via attention-weight manipulation. For detrimental head identification, we present two alternative strategies: AAT-GA that suited to limited floating-point compute (e.g., INT8/INT16 edge devices), and AAT-BP which offers greater optimization efficiency when resources allow. Experimental results demonstrate that AAT improves CLIP-family models across diverse downstream tasks, achieves far greater parameter efficiency than PEFT, and corroborates recent interpretability findings. We hope this study contributes to deeper understanding of CLIP and paves the way to efficiency of VLMs.
[…]
Supplementary Material
7 Benchmark Details and Evaluation Metrics
7.1 MS COCO
@@ 7.4 COCO-CN
7.5 Flickr30k-CNA
Flickr30k-CNA is a cleaned Chinese-translated version of Flickr30k, with captions generated by professional English–Chinese linguists. We process and evaluate it following the same protocol used for Flickr30k.
7.6 ImageNet-1k
@@
7.7 Cola
Cola [ray2024cola] is a compositional text-to-image retrieval benchmark designed to evaluate a model’s ability to localize and compose objects with their corresponding attributes. It includes both real-world images and synthetic 3D-rendered scenes. The task requires retrieving images that match the correct configuration of objects and attributes while rejecting distractors that contain the right components in incorrect arrangements. Cola consists of approximately 1.2k composed queries involving 168 objects and 197 attributes, distributed across roughly 30K images. It features two types of queries: single-object compositions (sourced from GQA, CLEVR, and PACO) and multi-object queries. Following the original paper, mean average precision (mAP) is used for evaluating single-object subsets, while mean accuracy is used for the multi-object subset.
@@ 7.7 Cola
8 Impact of Individual Attention Heads
[…]
9 Analysis of the Ablated Heads
9.1 Layer Distribution
@@
01234567891011055101033775544333303344556611224444333322553311447733layersnumCLIPChinese-CLIP    (a) ViT-B-based AAT 0246810121416182022055101015155588664455554455557755111144556677550121210108866669977665533338855555510106688666655776666776610106644layersnumCLIPChinese-CLIP     (b) ViT-L-based AAT 024681012141618202224262830055101015157710108888555544995577552277551188997777779977448888335544995599881010101066991212558844449977661188334488555566667788886644667744448866layersnumCLIPChinese-CLIP        (c) ViT-H-based AAT
@@ 9.2 Statistics
10 Computational Cost Analysis
We optimize AAT-GA by eliminating redundant overhead, as described in Section 5.5 of the paper. Detailed performance and runtime comparisons under different hyper-parameter configurations are provided in Tables 16 and 17.
@@ 10 Computational Cost Analysis
Table 16: Mean-R (mR) across CLIP model sizes for text-to-image retrieval on the MS COCO test set, along with the corresponding GA optimization runtime (T, in minutes). For clarity, “gen” refers to the number of generations, and “psize” denotes the population size.
@@
11 Evaluation on Compositional Retrieval
@@
12 Comparison with SFT
To further validate the effectiveness and practicality of AAT, we compare it against conventional supervised finetuning (SFT). We conduct this comparison using Chinese-CLIP, where AAT demonstrates more pronounced improvements than on the English counterpart, making performance differences easier to observe. Since AAT are optimized using 1k image-text pairs from COCO-CN, we train the vanilla Chinese-CLIP on the same dataset using the contrastive supervision objective employed in CLIP, following the recommended hyper-parameters from the OpenCLIP codebase [ilharco]. Specifically, we use a base learning rate of 5e-5 with a cosine decay schedule, the AdamW optimizer, a batch size of 256, and a weight decay of 1e-3. In experiments, we vary the number of training epochs from 1 to 64 to explore optimal performance across different model sizes. We evaluate the finetuned models on both the in-domain COCO-CN test set and the cross-domain Flickr30k-CNA test set for text-to-image retrieval. The results are presented in Figure 5.
SFT exhibits varying mean-R on COCO-CN across model sizes. It slightly outperforms AAT on ViT-B (++0.2%), achieves a larger gain on ViT-L (++2.2%), and performs comparably on ViT-H (-0.1%). However, its results are highly sensitive to the number of training epochs, likely due to overfitting or convergence to local minima under limited data. This instability is consistent across repeated trials, highlighting the challenges of SFT in data-scarce settings.
@@ 12 Comparison with SFT
13 Visualization
13.1 AAT Optimization Data
For the convenience of readers, we provide a visualization of the validation set $\mathcal{D}$ used for AAT optimization in Figure 6.
13.2 Failure Cases
As discussed in Appendix 11 and Section 5.2.1 of the paper, AAT fails on Cola-CLEVR and three non-natural subsets from ReCoS. Representative examples from these tasks are shown in Figure 7 and Figure 8, respectively. As illustrated, there exists a significant domain gap between the test images and those in $\mathcal{D}$. Moreover, certain tasks require models to perform high-level comprehension and reasoning, which exceeds the representational capacity of CLIP.
14 Limitations
@@
Figure 8: Visualization of non-natural images from three ReCoS-v1 subsets: code reasoning (CR), numerical calculation (NC), and text translation ($T$).